find the periodic withdrawals for the given annuity account.

$75,000 at 4%, paid out quarterly for 20 years.

Annuity question
Answer #1

PMT= Present value × (i/(1-(1+i)^-n))

n=mt = 4 quarters * 20 years=80 quarters

i= Interest rate /Quarters = 0.04/4= 0.01

Thus, PMT= 75000 × (0.01/ {1 - (1.01)^-80})

Thus PMT= periodic monthly intervals= 1366.41 $